# MLOps with Microservices:A Case Study on the Maritime Domain

### v1

@@ 9 Conclusions
  • alignment, related to coordinating four specialized teams working concurrently); and
@@
  • ubiquitous language, which created a shared vocabulary among all participants via the application of Domain Driven Design (DDD).
Contract-based design effectively mitigated coupling between components and improved alignment among the team, whereas the ubiquitous language enhanced communication with stakeholders and further strengthened alignment among the team. Therefore, this case study in the maritime domain illustrates how established software engineering practices can be successfully adapted to complex Machine Learning-Enabled Systems.

### v2

@@ 9 Conclusions
  • alignment, related to coordinating four specialized teams working concurrently; and
@@
  • ubiquitous language, which created a shared vocabulary among all participants via the application of Domain-Driven Design (DDD).
Contract-based design effectively mitigated coupling between components and improved alignment among the team, whereas the Ubiquitous Language enhanced communication with stakeholders and further strengthened alignment among the team. Therefore, this case study in the maritime domain illustrated how these software engineering techniques can be adapted for MLOps by extending their applicability to the data and model dimensions of complex Machine Learning–Enabled Systems.